### v2

@@ Abstract
Research in various fields is currently experiencing challenges regarding the reproducibility of results. This problem is also prevalent in machine learning (ML) research. The issue arises, for example, due to unpublished data and/or source code and the sensitivity of ML training conditions. Although different solutions have been proposed to address this issue, such as using ML platforms, the level of reproducibility in ML-driven research remains unsatisfactory. Therefore, in this article, we discuss the reproducibility of ML-driven research with three main aims: (i) identifying the barriers to reproducibility when applying ML in research as well as categorize the barriers to different types of reproducibility (description, code, data, and experiment reproducibility), (ii) discussing potential drivers such as tools, practices, and interventions that support ML reproducibility, as well as distinguish between technology-driven drivers, procedural drivers, and drivers related to awareness and education, and (iii) mapping the drivers to the barriers. With this work, we hope to provide insights and to contribute to the decision-making process regarding the adoption of different solutions to support ML reproducibility.
@@ 1 Introduction
As in many scientific fields [9, 68], e.g., medicine and behavior science, research in artificial intelligence (AI) in general, and machine learning (ML) in particular, faces crucial doubts over the reproducibility of research [36]. This has raised concerns about the reliability and validity of many scientific findings, risking diminishing confidence in the overall body of scientific knowledge. Misleading or irreproducible results can lead to wasted resources, hinder scientific progress, reduce trust in science, and impact decision-making in various fields [51, 29]. The “reproducibility crisis” has been a topic of discussion since the mid-2010s, particularly in the medical and behavioral science fields [9, 55]. Recently, this issue has also been associated with research involving ML methods [36]. Apart from the common challenges faced by many disciplines, such as limited sharing of data and source code, the use of ML introduces unique challenges for reproducibility, including sensitivity to ML training conditions, sources of randomness [65], and the increasing use of AutoML tools [42].
@@
While the literature on these issues is expanding [46, 30, 33, 40, 26, 4, 70, 27], there is still no comprehensive overview of the associated barriers and drivers. For example, in [27], the authors identify and categorize sources of irreproducibility in ML and how these sources affect conclusions drawn from ML experiments. However, this study does not investigate drivers to address these sources of irreproducibility. Thus, our paper provides a contextual categorization of the barriers and drivers to the four types of ML reproducibility proposed by (description, code, data, and experiment) [26], with specific reference to research in both Computer Science and Biomedical fields. We also propose a Drivers-Barriers-Matrix to summarize and visualize the results of the discussion. Such an analysis stands to clarify the current state regarding ML reproducibility, provide guidance for those new to the subject, and highlight potential gaps, solutions, and avenues for future exploration.
The paper is structured as follows: in Section 2, we clarify terms and working definitions. We then analyze the barriers to increased reproducibility of ML-driven research (Section 3), and next the drivers that support ML reproducibility, including different tools, practices, and interventions (Section 4). Finally, we map the barriers to the drivers to help determine the feasibility of various options for enhancing ML reproducibility (Section 5). We close the paper with a conclusion and an outlook into our future research in Section 6.
@@ 2 Defining Reproducibility
The concept of reproducibility can have different interpretations across various research fields and even within the same field [23]. To address this confusion, reproducibility is broadly defined and then further categorized into various types or degrees. A key distinction is made by Goodman et al. [25] who specify a fundamental division between whether we (i) mean reproducible in principle (termed “methods” reproducibility) due to sufficient description/sharing of methodologies, materials, etc., or (ii) whether results/conclusions actually prove to be reproducible when experiments or analyses are re-done. In the second category, they distinguish “results” and “inferential” reproducibility, depending on whether it is the analyses or inferences to broader conclusions that are reproduced.
Within ML research, widely accepted definitions that build further on these key distinctions have been proposed by  [30, 26]. We follow and build upon these latter definitions, and hence here outline them in some length.
Gundersen et al. [28] define reproducibility in general as “the ability of independent investigators to draw the same conclusions from an experiment by following the documentation shared by the original investigators”.
Relating to point (ii) of Goodman et al.’s schema, Gundersen and colleagues  [30, 26] further distinguish the targets of reproducibility, i.e., how closely an experiment can be reproduced:
@@
  • Analysis reproducibility does not require the reproduced experiment to have the same/similar outcome; however, if the same/similar analysis and, therefore also, interpretation can be made, an experiment is analysis reproducible.
  • Interpretation reproducibility does not require the reproduced experiment to have the same/similar outcome nor analysis but requires the interpretation to be the same as the original one. In such a case, an experiment is interpretation reproducible.
@@
This categorization aims to overcome the problem of ambiguity when making specific claims about the reproducibility of an experiment. Often in literature, authors write about reproducing the same “results” of an experiment. It is not apparent, however, in which cases they mean to achieve the same computational outcome, i.e., outputs of the algorithms, or whether they mean to reach the same analysis or interpretation. Therefore, achieving interpretation reproducibility is a more general and often less stringent requirement than achieving outcome reproducibility. This categorization is not specific to ML, but generally applicable to any research field that conducts computational experiments.
In addition, relating to point (i) of the Goodman et al. schema (“methods” reproducibility), [26] specifies reproducibility types to which methods can be made transparent through description or sharing. These four types are defined as R1 Description, R2 Code, R3 Data, and R4 Experiment. The lower the level of reproducibility, the less shared information there is, making the experiment more difficult to reproduce. Giving an example: In general, all published research experiments are accompanied by a textual description of the experiment. If this textual description is the only information shared by the authors, the research is categorized as R1 Description, which, according to this scheme, is the lowest type of reproducibility. In contrast, if all three building blocks, i.e., text, code, and data, are shared, the experiment can be categorized as R4 Experiment, which is the highest type of reproducibility. Furthermore, the distinction between R2 Code and R3 Data is defined by the textual description of the original experiment that is accompanied by either code or data, respectively.
Also, the different types of ML reproducibility exhibit an interplay between generalizability and reproducibility. R1 Description leads to the highest generalizability, but to the weakest reproducibility, while R4 Experiment leads to the highest reproducibility and lowest generalizability.
This relationship is illustrated in Figure 1. Furthermore, Figure 1 visualizes the building blocks associated with the different types of reproducibility. In what follows, unless stated otherwise, we use the definitions proposed by Gundersen and colleagues [26].
@@ 3 Barriers in ML Reproducibility
In the following, we discuss nine barriers to ML reproducibility, categorized into the four types of reproducibility mentioned in the previous section. Where applicable, we give examples within the research fields of Biomedical Science, and Computer Science.
@@ 3.1 R1 Description
    The evaluation metrics used to report results are not properly specified. There are many metrics, which can be used to evaluate ML models, e.g., accuracy, receiver-operator-curve (ROC), or mean-squared-error (MSE). It is important to define these metrics and also explain why they were used.
@@
    Often results are selectively reported, e.g., researchers may only provide results for the best performance out of many test runs, instead of properly assessing and reporting the performance over multiple runs. For example, in research on Natural Language Processing (NLP), often, reproductions turn out to perform worse than stated in the original publications [11].
Generally, it is important for studies to use a robust methodology and provide detailed reports, for other researchers to verify results and understand how analyses were conducted. While ML models have proven highly effective in Biomedical fields, studies often fall short in providing comprehensive and high-quality reporting. For example, in studies on predicting cardiometabolic risk from dietary patterns [56] or supporting the clinical management of diabetes [39], ML models were observed to be very promising. While this further enhances the promise of applying ML models for various clinical prediction tasks, there is a clear need for thorough reporting and validation of these models to allow for their integration into routine clinical care [39]. This is also true for the application of ML models for cancer imaging, where ML models often surpass radiologists in performance, but publications on these models lack the documentation details needed to reproduce the results [61].
@@
Another flaw, which has been observed in ML-based research and negatively affects reproducibility, is “spin”. It refers to the misuse of language to “intentionally or unintentionally affect the interpretation of study findings”. Accordingly, it is also understood as an inconsistency between the study results and the conclusions, in the sense that results are over-generalized or the claimed conclusions are not supported by the scientific method. This has been shown to impact both the interpretations and decision-making by readers [5]. In ML-based Biomedical research, the most common practice of spin includes recommending models for various applications without providing external validation in the same study. More concretely, the recommendation to use a model either in a clinical setting or for a different population is only validated in approximately 15% of the cases. Other observed instances of spin are invalid comparisons of results to previous studies, and the use of leading words and strong statements to make the results sound more prolific [5]. The use of spin in numerous publications might be partly attributed to the reward system in academia, which often values spin practices in research, as more prolific results are more likely to lead to accepted publications. Publication bias in general is a big problem in a variety of different regards. It seems that research from high-profile institutions is published more easily and subject to less scrutiny in review, and there also seems to be a tendency to prefer more complex solutions over simpler ones addressing the same problem [75].
@@ 3.2 R2 Code
Published ML research is often not accompanied by available data and code. Only one-third of researchers share data, and even fewer share their source code [37]. This can be for many reasons, such as the increasing pressure on researchers to publish quickly, which in turn does not allow them to polish the code and decreases the willingness to release the code, or simply due to privacy constraints with respect to sharing code. Sharing ML code correctly requires publishing seven pieces of information: hypothesis, prediction method, source code, hardware specifications, software dependencies, experiment setup, and experiment source code. Only if the ML code is accompanied by all these artifacts, it is specified sufficiently to allow reproducing an experiment [29]. Unfortunately, current research often does not meet this requirement, leading to reproducibility issues due to different software versions, hyperparameter settings, or hardware differences  [34, 11]. For example, in Computer Science, the lack of shared code and data significantly contributes to a lack of reproducibility [19]. Even when code and data are shared, they are often poorly documented, or the code is only provided as a skeleton instead of a fully executable code, which does not ensure reproducibility. Thus, shared code should also include the code and documentation for data preprocessing, hyperparameter tuning, and managing random seeds, as well as code for comparisons against baseline models.
@@ 3.3 R3 Data
The main barrier concerning R3 Data is that data is often not shared. This could be due to privacy concerns or a lack of incentives and motivation. Moreover, many benchmark and training datasets encounter challenges related to copyright, licensing, and longevity. These datasets may also raise ethical concerns, such as the unintentional inclusion of privacy-sensitive or harmful content, making it difficult to share the data for ML model training [57]. Similar to the sharing of source code, sharing only the datasets is insufficient. For proper use, it is also important to share specific splits, i.e., the training dataset, validation dataset, and test dataset [29]. Furthermore, data sharing needs to be accompanied by documentation specifying details about the provenance and preprocessing of data. This is crucial to investigate methodological errors such as data leakage and bias, which are described in the following.
@@
In practice, methodological issues such as data leakage often hinder the reproducibility of ML-based research [40]. This is due to the growing number of non-experts employing machine learning across different research fields [24], which is fuelled by the ease of application of ML libraries and no-code off-the-shelf AI tools. In essence, data leakage happens when data, on which the ML model should not be trained, leaks into the training process. Data leakage can be categorized into 3 subcategories [40]:
@@
    Use of non-legitimate data. For example, when the use of antihypertensive drugs is used to predict hypertension. This data is non-legitimate, since it would not be available in a real-world scenario and cannot realistically be used to predict hypertension for a new patient.
@@
Bias in ML refers to the error introduced by approximating a real-world problem, which may be complex, by a simplified model, often leading to systematic deviations from the real world. Furthermore, biases can arise when the model contains imbalances or reflects existing societal biases [47]. ML models, which are subject to bias, are prone to generalization issues, and are therefore potentially problematic for ML reproducibility. There are eight bias variants that can arise during the data handling phase of ML development [67]: (i) selection bias - using data not being representative of the target group, (ii) exclusion bias - excluding particular data samples based on the belief that they are unimportant, (iii) measurement bias - favoring certain measurement results, (iv) recall bias - labeling similar data samples differently, (v) survey bias - introducing data issues stemming from data collection surveys, (vi) confirmation bias - favoring information, which confirms previous beliefs, (vii) prejudice bias - including human-related prejudices in training data, and (viii) algorithmic bias - replicating or amplifying biases by the inner workings of the algorithms. Bias, such as selection bias, often leads to the issue of validity shrinkage in Biomedical science research [38]. For example, in obesity and nutritional research, ML is used to predict obesity, heart rate, or the risk of a heart attack, based on data of an individual. Here, validity shrinkage refers to the issue that a predictive model trained on a subset of data will most probably not perform well on new samples. The difference between predictive performance on known data and new data, however, is most often not accounted for in nutritional science, and therefore also leads to performance claims that cannot be reproduced [38].

### v3

@@ Abstract
Many research fields are currently reckoning with issues of poor levels of reproducibility. Some label it a “crisis”, and research employing or building Machine Learning (ML) models is no exception. Issues including lack of transparency, data or code, poor adherence to standards, and the sensitivity of ML training conditions mean that many papers are not even reproducible in principle. Where they are, though, reproducibility experiments have found worryingly low degrees of similarity with original results. Despite previous appeals from ML researchers on this topic and various initiatives from conference reproducibility tracks to the ACM’s new Emerging Interest Group on Reproducibility and Replicability, we contend that the general community continues to take this issue too lightly. Poor reproducibility threatens trust in and integrity of research results. Therefore, in this article, we lay out a new perspective on the key barriers and drivers (both procedural and technical) to increased reproducibility at various levels (methods, code, data, and experiments). We then map the drivers to the barriers to give concrete advice for strategies for researchers to mitigate reproducibility issues in their own work, to lay out key areas where further research is needed in specific areas, and to further ignite discussion on the threat presented by these urgent issues.
@@ 1 Introduction
Trustworthy AI requires reproducibility [70]. Unreliable results risk hindering scientific progress by wasting resources, reducing trust, slowing discovery, and undermining the foundation for future research [84, 45]. However, many scientific fields currently face crucial questions over the reproducibility of research findings [10]. Concerns of a “reproducibility crisis” have been most prominently raised in biomedical  [40, 59, 60, 35] and social  [88, 27, 51] sciences, but research employing Artificial Intelligence (AI) in general, and Machine Learning (ML) in particular, is also under scrutiny [57]. ML is becoming ever more deeply integrated into research methods, not just in computer science but across disciplines [32, 89]. Indeed, recipients of the 2024 Nobel Prizes for both chemistry and physics included ML researchers. Hence, issues regarding the reproducibility of ML raise urgent concerns about the reliability and validity of findings not only for computer scientists but for large swathes of cutting-edge scientific research across disciplines.
The causes of poor reproducibility can be technical, methodological, or cultural. Viewed from a high level, some causes, such as lack of sharing data and code, lack of or poor adherence to standards, suboptimal research design, or poor incentives, may be seen as common to many domains. Apart from the common challenges faced by other disciplines, the use of ML introduces unique obstacles for reproducibility, including sensitivity to ML training conditions, sources of randomness [101], inherent nondeterminism, costs (economic and environmental) of computational resources, and the increasing use of Automated-ML (AutoML) tools [67, 48]. Among the methodological and cultural aspects, specificities of ML research, like “data leakage”, as well as ML-specific issues regarding unobserved bias, lack of transparency, selective reporting of findings, and publishing cultures, each play a role as well. Indeed, this cultural aspect must not be underestimated. The culture of “publish or perish” pervades academia, pushing researchers to publish as many papers in the highest-ranked or most prestigious journals or conferences as possible [95]. In turn, this culture distorts incentives towards corner-cutting, giving rise to so-called “questionable research practices” and “design, analytic, or reporting practices that have been questioned because of the potential for the practice to be employed with the purpose of presenting biased evidence in favor of an assertion” [11].
@@
This paper aims to provide a detailed overview of reproducibility and its associated barriers and drivers in ML. This is urgently needed since, despite previous appeals from ML researchers on this topic, various initiatives from conference reproducibility tracks to the ACM’s new Emerging Interest Group on Reproducibility and Replicability, and an expanding literature on the topic [78, 46, 53, 64, 43, 5, 108, 44], we contend that the general community continues to take this issue too lightly. In addition, despite the growing literature, no such comprehensive overview exists. For example, in [44], the authors identify and categorize sources of irreproducibility in ML and how these sources affect conclusions drawn from ML experiments. However, this study does not investigate the drivers to address these sources of irreproducibility. Thus, our paper provides a contextual categorization of the barriers and drivers to the four types of ML reproducibility (description, code, data, and experiment) proposed by  [43], with specific reference to research in both computer science and biomedical fields. We also propose a Drivers-Barriers-Matrix to summarize and visualize the results of the discussion. Such an analysis stands to clarify the current state regarding ML reproducibility, to give concrete advice for strategies for researchers to mitigate reproducibility issues in their own work, to lay out key areas where further research is needed in specific areas, and to further ignite discussion on the threat presented by these urgent issues. The paper is structured as follows: in Section 2, we clarify terms and working definitions. We then analyze the barriers to increased reproducibility of ML-driven research (Section 3), and next, the drivers that support ML reproducibility, including different tools, practices, and interventions (Section 4). Here, we also provide a comparison of the strengths and potential limitations of these drivers. Finally, we map the barriers to the drivers to help determine the feasibility of various options for enhancing ML reproducibility (Section 5). We close the paper with a conclusion and an outlook into our future research in Section 6.
@@ 2 Defining Reproducibility
The concept of reproducibility can have different interpretations across various research fields and even within the same field [39]. To avoid confusion, we first specify our terms, broadly defining reproducibility and then further categorizing it into various types and degrees. The first distinction comes from Goodman et al. [42], who specify a fundamental division between whether we (i) mean reproducible in principle (termed “methods” reproducibility) due to sufficient description/sharing of methodologies, materials, etc., or (ii) whether results/conclusions actually prove to be reproducible when experiments or analyses are re-done. In the second category, they distinguish “results” and “inferential” reproducibility, depending on whether the analyses or inferences to broader conclusions are reproduced.
Within ML research, widely accepted definitions that build further on these key distinctions have been proposed by Gundersen et al. [46, 43]. We follow and build upon these latter definitions, and hence, here outline them at some length. Gundersen et al. [44] define reproducibility in general as “the ability of independent investigators to draw the same conclusions from an experiment by following the documentation shared by the original investigators”. Relating to point (ii) of Goodman et al.’s schema, Gundersen and colleagues  [46, 43] further distinguish the targets of reproducibility, i.e., how closely an experiment can be reproduced:
@@
  • Analysis reproducibility does not require the reproduced experiment to have the same/similar outcome; however, if the same/similar analysis and, therefore, also interpretation can be made, an experiment is analysis reproducible.
  • Interpretation reproducibility does not require the reproduced experiment to have the same/similar outcome nor analysis but requires the interpretation to be the same as the original one.
This categorization aims to overcome the problem of ambiguity when making specific claims about the reproducibility of an experiment. Often in literature, authors write about reproducing the same “results” of an experiment. It is not apparent, however, in which cases they mean to achieve the same computational outcome, i.e., outputs of the algorithms, or whether they mean to reach the same analysis or interpretation. Therefore, achieving interpretation reproducibility is a more general and often less stringent requirement than achieving outcome reproducibility. This categorization is not specific to ML, but is generally applicable to any research field that conducts data analysis and interpretation.
In addition, relating to point (i) of the Goodman et al. schema (“methods” reproducibility), [43] specifies reproducibility types to which methods can be made transparent through description or sharing. These four types are defined as R1 Description, R2 Code, R3 Data, and R4 Experiment. The lower the level of reproducibility, the less shared information is shared, making the study more difficult to reproduce. As an example, in general, all published research experiments are accompanied by a textual description of the experiment. If this textual description is the only information shared by the authors, the research is categorized as R1 Description, which, according to this scheme, is the minimal kind of reproducibility. In contrast, if all three building blocks, i.e., text, code, and data, are shared, the experiment can be categorized as R4 Experiment, the most expansive kind of reproducibility. Furthermore, the distinction between R2 Code and R3 Data is defined by whether the textual description is accompanied by either code or data, respectively.
@@
Also, the different types of ML reproducibility exhibit an interplay between generalizability and reproducibility. R1 Description leads to strong generalizability but to weak reproducibility, while R4 Experiment leads to stronger reproducibility but weaker generalizability. What this means in sum is that rerunning the same code on the same data using the same description will make it likelier to obtain the same results, but those results might still be wrong due to errors or biases in any of those elements. On the other hand, building code from scratch and using alternative datasets for analysis will show that the techniques give similar results across contexts and, hence, higher levels of confidence in the generalizability of findings. These relationships are illustrated in Figure 1. In what follows, unless stated otherwise, we use the definitions proposed by Gundersen and colleagues [43].
@@ 3 Barriers in ML Reproducibility
Next, we discuss nine barriers to ML reproducibility, categorized into the four types of reproducibility mentioned beforehand. Where applicable, we give examples within the research fields of biomedical science and computer science.
@@ 3.1 R1 Description
    The evaluation metrics used to report results are not properly specified. There are many metrics which can be used to evaluate ML models, e.g., accuracy, receiver-operator-curve (ROC), or mean-squared-error (MSE). It is important to define these metrics and also explain why they were used.
@@
    Often results are selectively reported, e.g., researchers may only provide results for the best test run out of many test runs, instead of properly assessing and reporting average values and variances [14].
Generally, it is important for studies to use a robust methodology and provide detailed reports so that other researchers can verify results and understand how analyses were conducted. While ML models have proven highly effective in biomedical fields, studies often fall short in providing comprehensive and high-quality reporting. For example, in studies on predicting cardiometabolic risk from dietary patterns [90] or supporting the clinical management of diabetes [63], ML models were observed to be very promising. While this further enhances the promise of applying ML models for various clinical prediction tasks, there is a clear need for thorough reporting and validation of these models to allow for their integration into routine clinical care [63]. This is also true for the application of ML models for cancer imaging, where ML models often surpass radiologists in performance, but publications on these models lack the documentation details needed to reproduce the results [97].
@@
Another issue commonly observed in ML-based research that negatively affects reproducibility is “spin”. It refers to the misuse of language to “intentionally or unintentionally affect the interpretation of study findings”. It is also understood as an inconsistency between the study results and the conclusions, in the sense that results are over-generalized or the claimed conclusions are not supported by the scientific method. This has been shown to impact both the interpretations and decision-making by readers [6]. In ML-based biomedical research, the most common practice of spin includes recommending models for various applications without providing external validation in the same study. More concretely, the recommendation to use a model either in a clinical setting or for a different population is only validated in approximately 15% of the cases. Other observed instances of spin are invalid comparisons of results to previous studies and the use of leading words and strong statements to make the results sound more significant [6]. The prevalence of spin can perhaps be attributed in large part to the academic culture of “publish or perish” and its associated reward systems. Valuing and rewarding perceived novelty and potential impact over basic rigor and responsible reporting can lead researchers to inflate claims in hopes of acceptance in the most prestigious venues. It can also skew the literature in other ways, leading to so-called “publication bias” [105]. Here, in addition to spin and the aforementioned selective reporting of (usually positive) results, the role of the peer review system is also in question, given known biases on the part of reviewers that can lead to preferential treatment for researchers from specific regions, institutions or demographics, or for certain types of research [73]. Other kinds of bias, such as “complexity bias” (tendency to prefer complicated over simple results and explanations), are also known to influence acceptance decisions [115].
Finally, we note how a core aspect of computer science culture may exacerbate these issues, namely the importance of conference rankings. Within computer science, the prime mode of publication is within conference proceedings, with conferences ranked A* to C by bodies such as ICORE [^1] https://portal.core.edu.au/conf-ranks/. To date, surprisingly little has been said regarding the analogous nature of such conference rankings to other metrics like the Journal Impact Factor, where a rich literature exists critiquing its worth as an indicator of quality or impact for individual pieces of work [72]. Although elaboration at length on this issue is outside the scope of this article, we suggest this as an underexplored topic for future research. Such research can build on a rich evidence-base exploring downstream ill-effects of badly designed or misused metrics, including distortion of incentives [110], inviting manipulation or gaming [15], goal displacement and task reduction [102], and influencing core academic values [20].
@@ 3.2 R2 Code
Published ML research is often not accompanied by available data and code. Only one-third of researchers share data, and even fewer share their source code [58]. This can be attributed to several factors, such as the increasing pressure on researchers to publish quickly, which often leaves insufficient time to refine code and decreases the willingness to share it. Additionally, concerns about intellectual property may further discourage researchers from releasing their code. According to Gundersen and Kjensmo [45], sharing ML code to facilitate reproducibility requires publishing seven pieces of information: hypothesis, prediction method, source code, hardware specifications, software dependencies, experiment setup, and experiment source code. Unfortunately, current research rarely meets these requirements, leading to reproducibility issues due to different software versions, hyperparameter settings, or hardware differences [54, 14]. For example, research in recommender systems has struggled with the lack of shared code, significantly contributing to a lack of reproducibility. Even when code is shared, it is often incomplete, poorly documented, or limited to pseudocode or skeletal implementations rather than fully executable code [29]. To address this, shared code should encompass comprehensive documentation, including scripts for data preprocessing, hyperparameter tuning, management of random seeds, and implementations for comparisons against baseline models.
@@ 3.3 R3 Data
The main reproducibility barrier associated with R3 Data is that data is simply not shared or made publicly available most of the time [58]). A review in biomedical research, specifically radiomics, investigated 257 recent ML publications and found that only 16 of them shared data or used publicly accessible datasets [33]. This could be due to privacy concerns or a lack of incentives and motivation. Moreover, many benchmark and training datasets encounter challenges related to copyright, licensing, and longevity. These datasets may also raise ethical concerns, such as the unintentional inclusion of privacy-sensitive or harmful content, making it difficult to share the data for ML model training [91]. Similar to the sharing of source code, sharing only the datasets is insufficient without sufficiently detailed levels of documentation. For proper use, it is also important to share specific splits, i.e., the training dataset, validation dataset, and test dataset [45]. Furthermore, data sharing needs to be accompanied by documentation specifying details about the provenance and preprocessing of data. Significant recent initiatives aiming at improvement here are: Croissant, a unified format for machine learning datasets that integrates metadata, resource descriptions, data structure, and default ML semantics in a single file [^2] https://github.com/mlcommons/croissant, and MLCommons, which is working towards open benchmarks and public data [^3] https://github.com/mlcommons. In addition, RO-Crate is a standard for packaging data and other research objects together with data to enable reuse and reproducibility [^4] https://www.researchobject.org/ro-crate/.
Standardising and mainstreaming these practices is essential for validation and checking of methods. We next discuss two common methodological errors related to data, data leakage and bias, and their impact on reproducibility.
@@
In practice, methodological issues such as data leakage (also referred to as target leakage) often hinder the reproducibility of ML-based research [64]. This is due to the growing number of non-experts employing machine learning across different research fields [41], which is fueled by the ease of application of ML libraries and no-code off-the-shelf AI tools. In essence, data leakage happens when data on which the ML model should not be trained leaks into the training process. Data leakage can be categorized into 3 subcategories [64]:
@@
    Use of non-legitimate data. For example, when the use of antihypertensive drugs is used as a feature to predict hypertension. This data is non-legitimate, since it would not be available in a real-world scenario (people are prescribed those drugs because of a hypertension diagnosis) and would be useless for predicting hypertension in undiagnosed patients.
@@
Bias in ML refers to the error introduced by approximating a real-world problem, which may be complex, by a simplified model, often leading to systematic deviations from the real world. Furthermore, biases can arise when the model contains imbalances or reflects existing societal biases [79]. ML models, which are subject to bias, are prone to generalization issues, and are therefore potentially problematic for ML reproducibility. There are eight kinds of bias that can arise during the data handling phase of ML development [104]: (i) selection bias - using data not being representative of the target group, (ii) exclusion bias - excluding particular data samples based on the belief that they are unimportant, (iii) measurement bias - favoring certain measurement results, (iv) recall bias - labeling similar data samples differently, (v) survey bias - introducing data issues stemming from data collection surveys, (vi) confirmation bias - favoring information, which confirms previous beliefs, (vii) prejudice bias - including human-related prejudices in training data, and (viii) algorithmic bias - replicating or amplifying biases by the inner workings of the algorithms. Bias, such as selection bias, often leads to the issue of validity shrinkage in biomedical science research [61]. For example, in obesity and nutritional research, ML is used to predict obesity, heart rate, or the risk of a heart attack based on data from an individual. Here, validity shrinkage refers to the issue that a predictive model trained on a subset of data will most probably not perform well on new samples. The difference between predictive performance on known data and new data, however, is most often not accounted for in nutritional science, and therefore also leads to performance claims that cannot be reproduced [61].